OpenStudy (anonymous):

how do i find b/a in the following : b/5=9/a=3? thanks

OpenStudy (anonymous):

\[\frac{b}{5} = \frac{9}{a} = 3\]

OpenStudy (anonymous):

You can find \(a\) and \(b\) and then take the ratio.. :)

OpenStudy (anonymous):

\[\frac{b}{5} = 3 \quad \quad \frac{9}{a} = 3\] Find \(a\) and \(b\)..

OpenStudy (anonymous):

a=3 and b=15?

OpenStudy (anonymous):

oh then its becomes 5 got it

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

But wait..

OpenStudy (anonymous):

Yes, finally you will get \(5\)..

OpenStudy (anonymous):

so a=1 and b=5